# Optimal streaming algorithm for detecting $\ell_{2}$ heavy hitters in random order streams

### v1

@@ Abstract
Given a stream $x_{1},x_{2},\dots,x_{n}$ of items from a Universe $U$ of size $\mathrm{poly}(n)$, and a parameter $\epsilon>0$, an item $i\in U$ is said to be an $\ell_{2}$ heavy hitter if its frequency $f_{i}$ in the stream is at least $\sqrt{\epsilon F_{2}}$, where $F_{2}=\sqrt{\sum_{i\in U}f_{i}^{2}}$. Efficiently detecting such heavy hitters is a fundamental problem in data streams and has several applications in both theory and in practice. The classical CountSketch algorithm due to Charikar, Chen, and Farach-Colton [2004], was the first algorithm to detect $\ell_{2}$ heavy hitters using $O\left(\frac{\log^{2}n}{\epsilon}\right)$ bits of space, and their algorithm is optimal for streams with deletions. A follow-up paper due to Braverman, Chestnut, Ivkin, and Woodruff [2016], gave the CountSieve algorithm for insertion-only streams, with an improved space bound. Their algorithm requires only $O\left(\frac{\log(1/\epsilon)}{\epsilon}\log n\log\log n\right)$ space. Note that any algorithm requires at least $O\left(\frac{1}{\epsilon}\log n\right)$ space to output $O(1/\epsilon)$ heavy hitters in the worst case. So for constant $\epsilon$, the space usage of the CountSieve algorithm is asymptotically only a $\log\log n$ factor worse than the optimum bound. A later work due to Braverman, Chestnut, Ivkin, Nelson, Wang, and Woodruff [2017] gave the BPTree algorithm which gets rid of this additional $\log\log n$ factor and detects $\ell_{2}$ heavy hitters in insertion-only streams using only $O\left(\frac{\log(1/\epsilon)}{\epsilon}\log n\right)$ space. While their algorithm achieves optimal space bound for constant $\epsilon$, their bound could be sub-optimal for $\epsilon=o(1)$. In this work, we show that for random order streams, where the stream elements can be adversarial but their order of arrival is uniformly random, it is possible to achieve the optimal space bound of $O\left(\frac{1}{\epsilon}\log n\right)$ for every $\epsilon=\Omega\left(\frac{1}{2^{\sqrt{\log n}}}\right)$. We also show that for partially random order streams where only the heavy hitters are required to be uniformly distributed in the stream, it is possible to achieve the same space bound, but with an additional assumption that the algorithm is given a constant approximation to $F_{2}$ in advance.
We mainly exploit the fact that if the stream is divided into windows of length roughly $\frac{n}{\sqrt{\epsilon F_{2}}}$, with high probability, for every heavy hitter, every $\log n$ consecutive windows has an occurrence of that heavy hitter, and develop a hierarchical sample-and-check framework that efficiently detects them. The window size is crucial to the algorithm. For random order streams, we show that it is possible to approximately estimate the window size after the algorithm sees a small constant fraction of the stream.
@@ 1 Introduction
Another well-studied guarantee is the $\ell_{2}$ guarantee, where given a parameter $\epsilon\in(0,1]$, the algorithm has to recover every element $i\in U$ such that $f_{i}^{2}\geq\epsilon\|f\|_{2}^{2}$ and not output any element $i\in U$ such that $f_{i}^{2}<\delta\|f\|_{2}^{2}$, where $\|f\|_{2}$ denotes the norm of the frequency vector of the elements in the stream. Note that this is a stronger guarantee compared to the $\ell_{1}$ guarantee. For example, consider a stream where $\|f\|_{2}^{2}=\Theta(n)$. For a constant $\epsilon$, the $\ell_{1}$ guarantee asks only to output elements that occupy a constant fraction of positions in the stream, while the $\ell_{2}$ guarantee asks for any element that has frequency $\Omega(\sqrt{n})$, which occupies only a negligible portion of the stream. The $\mathsf{CountSketch}$ algorithm due to Charikar, Chen, and Farach-Colton can recover heavy hitters under the $\l _{2}$ guarantee, with at least a constant probability, using $O(\log^{2}n/\epsilon)$ bits of space [18]. A follow-up work due to Jowhari, Sağlam, and Tardos shows that $\mathsf{CountSketch}$ can be extended to output heavy hitters under the $\ell_{p}$ guarantee for every $p\leq 2$ using space at most $O(\log^{2}n/\epsilon)$ and that the algorithm is optimal for every $p$ in this range, in the turnstile setting [17]. An earlier work due to Do ba, Indyk, Price, and Woodruff [19] proved this result for $p=2$. While $\mathsf{CountSketch}$ is optimal for turnstile streams, it is natural to ask if there is a more space-efficient algorithm in the insertion-only model, similar to the $\ell_{1}$ case. This question was first answered positively by Braverman, Chestnut, Ivkin, and Woodruff who gave the $\mathsf{CountSieve}$ algorithm which requires only $O\left(\frac{1}{\epsilon}\log\left(\frac{1}{\epsilon}\right)\log n\log\log n\right)$ space to output heavy hitters under the $\ell_{2}$ guarantee, with at least a constant success probability [20]. The same authors, along with Nelson and Wang, later gave an improved algorithm called $\mathsf{BPTree}$ which uses only $O\left(\frac{1}{\epsilon}\log\left(\frac{1}{\epsilon}\right)\log n\right)$ space to achieve a constant success probability [21]. The primary goal of this paper is to further reduce their space bound and obtain an optimal $O(\log n/\epsilon)$ space algorithm and we achieve this for random order streams. In addition, we also show that for partially random order streams when the heavy hitters are uniformly distributed across the stream (while the non-heavy elements could be adversarially positioned), it is possible to recover the $l_{2}$ heavy hitters, assuming we know the $l_{2}$ norm of the frequency vector in advance.
1.1 Our results
We now formally state our main results. We first begin with some definitions.
Definition 1.1 ($\ell_{2}$ norm).
The $\ell_{2}$ norm of a stream $S$, denoted by $\|S\|_{2}$, is defined to be the $\ell_{2}$ norm of the frequency vector $\mathbf{f}=(f_{i})_{i\in U}$ where $f_{i}$ is the number of occurrences of element $i$ in the stream.
Definition 1.2 ($\ell_{2}$ $\epsilon$-heavy hitter).
For $0<\epsilon\leq 1$, we say that $x$ is an $\epsilon$-heavy hitter if $f_{x}^{2}\geq\epsilon\|S\|_{2}$.
Random order streams are well-studied in the streaming literature, and understanding the space complexity in this model is an interesting problem of its own [22, 23, 24, 25]. The difference between random order streams and adversarial streams is that in random order streams, the elements in the stream are presented in a uniformly random order to the streaming algorithm, and the algorithm has to do well on most orders. Below, we state our result for detecting $\ell_{2}$ heavy hitters in random order streams.
Theorem 1.3 (Random order streams).
$\exists n_{0},c_{0}\in\mathbb{N}$ such that for every $n\geq n_{0}$ and $\epsilon\geq c_{0}\left(\frac{1}{2^{\sqrt{\log n}}}\right)$, there is a single-pass streaming algorithm that with probability at least $9/10$, finds every $\ell_{2}$ $\epsilon$-heavy hitter and reports no element that is not an $\ell_{2}$ $\epsilon/256$-heavy hitter, in random order streams of length $n$, using $O(\log n/\epsilon)$ bits of space.
We also extend our results to partially random order streams, where the positions of the heavy hitters are random, while the non-heavy elements can be adversarially positioned (see Definition 3.3 for a formal definition). However, in this case, we need an additional assumption that the $\ell_{2}$ norm of the stream is known in advance to the algorithm.
@@ 1.1 Our results
$\exists n_{0},c_{0}\in\mathbb{N}$ such that for every $n\geq n_{0}$ and $\epsilon\geq c_{0}\left(\frac{1}{2^{\sqrt{\log n}}}\right)$, there is a single-pass streaming algorithm that with probability at least $9/10$, finds every $\ell_{2}$ $\epsilon$-heavy hitter and reports no element that is not an $\ell_{2}$ $\epsilon/256$-heavy hitter, in partially random order streams of length $n$, using $O(\log n/\epsilon)$ bits of space, assuming that the $\ell_{2}$ norm of the stream is known in advance to the algorithm.
The above result should be viewed as a first step towards generalizing our result beyond random order streams. Assuming that the heavy hitters are uniformly distributed, while the non-heavy elements are placed by an adversary, is a natural starting point. For example, the heavy hitters maybe generated by a stochastic process, while the non-heavy elements could be adversarial noise. Such partially random ordered streams have been explored in the literature (see, for example, [26]) and serve as an important bridge between the random order and the adversarial streaming model. Our other assumption that the streaming algorithm needs to know the $\ell_{2}$ norm of the stream in advance is not as natural as our first assumption, and we state it this way purely for technical reasons. While one could imagine using an $\ell_{2}$ norm tracker algorithm like in the $\mathsf{BPTree}$ algorithm to get rid of this assumption, unfortunately if the stream order is not fully random, this could incur an additional factor of $\log(1/\epsilon)$ in space, and hence we would not gain any advantage in space compared to the $\mathsf{BPTree}$ algorithm. Therefore, we leave it as an interesting open problem whether one could prove Theorem 1.4 without the assumption about the $\ell_{2}$ norm of the stream, or if there is a lower bound in this case. We prove Theorems 1.3 and 1.4 at the end of Section 6.
@@ 2 Technical overview
2.1 Warmup
To motivate our new algorithm for super heavy hitter, let us first consider a special case of the stream as a warmup. Consider a stream of length $n$ consisting of a super heavy hitter $x^{*}$ with frequency $C\sqrt{n}$ for some large constant $C$, and other elements with frequency $1$. Dividing the stream into windows of length $W:=\sqrt{n}/C$, let us assume $x^{*}$ has one occurrence in every window. $x^{*}$ is super heavy in the stream, and we want to find $x^{*}$ with high probability using $O(\log n)$ bits of space.
@@ 2.3 $K$-super heavy hitter
The main challenge here is that we cannot remember in which window we sampled the element. Our solution first breaks the $\mathsf{check}$ into $O(\log n)$ iterations with geometrically increasing numbers of windows. In iteration $l$, we aim to verify $2^{l}$ consecutive windows. We always start the iteration from the next window $j$ such that $j$ is a multiple of $2^{l}$. [^3] This allows us to not remember when we started the iteration by only remembering the current window number. We can afford to remember $l$ (taking $O(\log l)=O(\log\log n)$ bits), but not an index within a $2^{l}$ range. The current window number takes $O(\log n)$ bits, but can be shared across all hash functions. In this case, from the fact that we are current executing iteration $l$, and the index of the current window, we will be able to derive a range which $i$ may be in. One can verify that this is a range $I$ of $2^{l}$ windows.
@@ 3.2 Concentration bounds
A theorem due to Dubhashi and Ranjan [28, Proposition 5] states that Chernoff-Hoeffding bounds can be applied on negatively associated random variables. In particular, we have the following.
@@
The following proposition from [28, Proposition 7] states some useful properties of negatively associated random variables.
@@
For $j\in[n]$, let $B_{j}$ be the indicator random variable for the event that ball numbered $j$ is sampled. Dubhashi, Priebe, and Ranjan proved in [29, Corollary 11] that the random variables $(B_{j}:j\in[n])$ are negatively associated. It follows from Item 1 in Proposition 3.12 that $(B_{j}:i+1\leq j\leq i+\ell)$ are also negatively associated. Since $X_{\ell,i}=\sum_{j=i+1}^{i+\ell}B_{j}$, $\mathbb{E}[X_{\ell,j}]=\ell k/n$, and the variables $B_{j}$ are identically distributed, the desired bound follows from applying Theorem 3.10. ∎
@@ 4 Streaming algorithm for detecting $\ell_{2}$ super heavy hitter
Let $\mathcal{S}$ be an input stream divided into $L$ windows, each of length $W$. Let $m=LW$ denote the total length of the stream. Let $U$ be a universe of size at most $\mathrm{poly}\,n$ [^4] Every element in the universe can be represented using $O(\log n)$ bits.. $K=(\log n)^{100}$ is a fixed parameter. We say that an element $x$ is a $K$-super heavy hitter if $f_{x}^{2}\geq K\cdot\sum_{y\neq x}f_{y}^{2}$, or equivalently it is a $\left(1-\frac{1}{K+1}\right)$-heavy hitter by Definition 1.2. We are given a subset of the universe $V\subset U$, and we focus on the substream consisting of elements in $V$. The positive case is that there exists a $K$-super heavy hitter $x^{*}\in V$ (with respect to the substream restricted to $V$) of frequency $L$ and evenly distributed in the original stream. In this section, we develop algorithms that return YES if we are in the positive case, with high probability, return NO if we are far from it (i.e., every element in $V$ has frequency at most $K^{-1/8}L$), and with further guarantees for the intermediate cases. In the following section, we first give a streaming algorithm for the case when $L\geq K^{3/16}$, and in the end, we handle the case when $L<K^{3/16}$ separately.
@@ 5 Streaming algorithm for detecting $\ell_{2}$ heavy hitter
[…]

### v2

# Optimally detecting uniformly-distributed $\ell_{2}$ heavy hitters in data streams

@@ Abstract
Given a stream $x_{1},x_{2},\dots,x_{n}$ of items from a Universe $U$ of size $\mathrm{poly}(n)$, and a parameter $\epsilon>0$, an item $i\in U$ is said to be an $\ell_{2}$ heavy hitter if its frequency $f_{i}$ in the stream is at least $\sqrt{\epsilon F_{2}}$, where $F_{2}={\sum_{i\in U}f_{i}^{2}}$. Efficiently detecting such heavy hitters is a fundamental problem in data streams and has several applications in both theory and in practice. The classical CountSketch algorithm due to Charikar, Chen, and Farach-Colton [2004], was the first algorithm to detect $\ell_{2}$ heavy hitters using $O\left(\frac{\log^{2}n}{\epsilon}\right)$ bits of space, and their algorithm is optimal for streams with deletions. A follow-up paper due to Braverman, Chestnut, Ivkin, and Woodruff [2016], gave the CountSieve algorithm for insertion-only streams, with an improved space bound. Their algorithm requires only $O\left(\frac{\log(1/\epsilon)}{\epsilon}\log n\log\log n\right)$ space. Note that any algorithm requires at least $\Omega\left(\frac{1}{\epsilon}\log n\right)$ space to output $O(1/\epsilon)$ heavy hitters in the worst case. So for constant $\epsilon$, the space usage of the CountSieve algorithm is asymptotically only a $\log\log n$ factor worse than the optimum bound. A later work due to Braverman, Chestnut, Ivkin, Nelson, Wang, and Woodruff [2017] gave the BPTree algorithm which gets rid of this additional $\log\log n$ factor and detects $\ell_{2}$ heavy hitters in insertion-only streams using only $O\left(\frac{\log(1/\epsilon)}{\epsilon}\log n\right)$ space. While their algorithm achieves optimal space bound for constant $\epsilon$, their bound could be sub-optimal for $\epsilon=o(1)$. For random order streams, where the stream elements can be adversarial but their order of arrival is uniformly random, Braverman, Garg, and Woodruff [2020] showed that it is possible to achieve the optimal space bound of $O\left(\frac{1}{\epsilon}\log n\right)$ for every $\epsilon=\Omega\left(\frac{1}{2^{\sqrt{\log n}}}\right)$.
In this work, we generalize their result to partially random order streams where only the heavy hitters are required to be uniformly distributed in the stream. We show that it is possible to achieve the same space bound, but with an additional assumption that the algorithm is given a constant approximation to $F_{2}$ in advance. We mainly exploit the fact that if the stream is divided into windows of length roughly $\frac{n}{\sqrt{\epsilon F_{2}}}$, with high probability, for every heavy hitter, every $\log n$ consecutive windows has an occurrence of that heavy hitter, and develop a hierarchical sample-and-check framework that efficiently detects them. The window size is crucial to the algorithm.
@@ 1 Introduction
Another well-studied guarantee is the $\ell_{2}$ guarantee, where given a parameter $\epsilon\in(0,1]$, the algorithm has to recover every element $i\in U$ such that $f_{i}^{2}\geq\epsilon\|f\|_{2}^{2}$ and not output any element $i\in U$ such that $f_{i}^{2}<\delta\|f\|_{2}^{2}$, where $\|f\|_{2}$ denotes the norm of the frequency vector of the elements in the stream. Note that this is a stronger guarantee compared to the $\ell_{1}$ guarantee. For example, consider a stream where $\|f\|_{2}^{2}=\Theta(n)$. For a constant $\epsilon$, the $\ell_{1}$ guarantee asks only to output elements that occupy a constant fraction of positions in the stream, while the $\ell_{2}$ guarantee asks for any element that has frequency $\Omega(\sqrt{n})$, which occupies only a negligible portion of the stream. The $\mathsf{CountSketch}$ algorithm due to Charikar, Chen, and Farach-Colton can recover heavy hitters under the $\l _{2}$ guarantee, with at least a constant probability, using $O(\log^{2}n/\epsilon)$ bits of space [9]. A follow-up work due to Jowhari, Sağlam, and Tardos shows that $\mathsf{CountSketch}$ can be extended to output heavy hitters under the $\ell_{p}$ guarantee for every $p\leq 2$ using space at most $O(\log^{2}n/\epsilon)$ and that the algorithm is optimal for every $p$ in this range, in the turnstile setting [19]. An earlier work due to Do ba, Indyk, Price, and Woodruff [1] proved this result for $p=2$. While $\mathsf{CountSketch}$ is optimal for turnstile streams, it is natural to ask if there is a more space-efficient algorithm in the insertion-only model, similar to the $\ell_{1}$ case. This question was first answered positively by Braverman, Chestnut, Ivkin, and Woodruff who gave the $\mathsf{CountSieve}$ algorithm which requires only $O\left(\frac{1}{\epsilon}\log\left(\frac{1}{\epsilon}\right)\log n\log\log n\right)$ space to output heavy hitters under the $\ell_{2}$ guarantee, with at least a constant success probability [7]. The same authors, along with Nelson and Wang, later gave an improved algorithm called $\mathsf{BPTree}$ which uses only $O\left(\frac{1}{\epsilon}\log\left(\frac{1}{\epsilon}\right)\log n\right)$ space to achieve a constant success probability [6]. For random order streams where the order of arrival of elements is uniformly random, Braverman, Garg, and Woodruff [5] gave an optimal algorithm that uses only $O(\log n/\epsilon)$ space and achieves constant success probability, for every $\epsilon=\Omega(1/2^{\sqrt{\log n}})$. The primary goal of this paper is to extend their result to more general settings. In particular, we show that for partially random order streams when the heavy hitters are uniformly distributed across the stream (while the non-heavy elements could be adversarially positioned), it is still possible to recover the $\epsilon$-heavy $l_{2}$ heavy hitters, using only $O(\log n/\epsilon)$ space with constant success probability, for every $\epsilon=\Omega(1/2^{\sqrt{\log n}})$, assuming have an estimate of the $l_{2}$ norm of the frequency vector and access to a random oracle.
1.1 Our result
We now formally state our main theorem.
@@ 1.1 Our result
$\exists n_{0},c_{0}\in\mathbb{N}$ such that for every $n\geq n_{0}$ and $\epsilon\geq c_{0}\left(\frac{1}{2^{\sqrt{\log n}}}\right)$, there is a single-pass streaming algorithm that with probability at least $9/10$, finds every $\ell_{2}$ $\epsilon$-heavy hitter and reports no element that is not an $\ell_{2}$ $\epsilon/256$-heavy hitter, in partially random order streams of length $n$, using $O(\log n/\epsilon)$ bits of space, assuming that the $\ell_{2}$ norm of the stream is known in advance to the algorithm and access to a random oracle.
The above result should be viewed as a first step towards generalizing [5] beyond random order streams. Assuming that the heavy hitters are uniformly distributed, while the non-heavy elements are placed by an adversary, is a natural starting point. For example, the heavy hitters maybe generated by a stochastic process, while the non-heavy elements could be adversarial noise. Such partially random ordered streams have been explored in the literature (see, for example, [10]) and serve as an important bridge between the random order and the adversarial streaming model. Our other assumption that the streaming algorithm needs to know the $\ell_{2}$ norm of the stream in advance is not as natural as our first assumption, and we state it this way purely for technical reasons. While one could imagine using an $\ell_{2}$ norm tracker algorithm like in the $\mathsf{BPTree}$ algorithm to get rid of this assumption, unfortunately if the stream order is not fully random, this could incur an additional factor of $\log(1/\epsilon)$ in space, and hence we would not gain any advantage in space compared to the $\mathsf{BPTree}$ algorithm. [^2] Such an algorithm that geometrically guesses $\ell_{2}$ must start the subroutine with the correct guess at the latest, say when there is still half of the stream left. Otherwise, there are not enough occurrences of the heavy hitters to successfully detect them. However, imagine a case where the true $\ell_{2}$-norm is $F=n^{1.5}$, and there is only one heavy hitter in the stream, which has squared frequency $\epsilon F$. All other (light) elements occur once in the first half. Then we can only rule out the $\ell_{2}$ smaller than $O(\epsilon F)$ only when we have processed the first half of the stream. Hence, we must maintain $O(\log(1/\epsilon))$ consecutive geometric guesses. Therefore, we leave it as an interesting open problem whether one could prove Theorem 1.1 without the assumption about the $\ell_{2}$ norm of the stream, or if there is a lower bound in this case. We prove Theorem 1.1 at the end of Section 6.
@@ 2 Technical overview
2.1 Warmup and summary of [5]
To motivate our new algorithm for super heavy hitter, let us first consider a special case of the stream as a warmup. Consider a stream of length $n$ consisting of a super heavy hitter $x^{*}$ with frequency $C\sqrt{n}$ for some large constant $C$, and other elements with frequency $1$. Dividing the stream into windows of length $W:=\sqrt{n}/C$, let us assume $x^{*}$ has one occurrence in every window. $x^{*}$ is super heavy in the stream, and we want to find $x^{*}$ with high probability using $O(\log n)$ bits of space. Note that such partially random order streams actually have random order. In fact, the algorithm presented below is a (simplification) of the prior work [5] for random-order streams. [^3] The prior work [5] uses a different setting of parameters, and their algorithm for this case succeeds with constant probability instead of with high probability. On the other hand, their algorithm does not assume free randomness.
@@ 2.3 $K$-super heavy hitter
The main challenge here is that we cannot remember in which window we sampled the element. Our solution first breaks the $\mathsf{check}$ into $O(\log n)$ iterations with geometrically increasing numbers of windows. In iteration $l$, we aim to verify $2^{l}$ consecutive windows. We always start the iteration from the next window $j$ such that $j$ is a multiple of $2^{l}$. [^5] This allows us to not remember when we started the iteration by only remembering the current window number. We can afford to remember $l$ (taking $O(\log l)=O(\log\log n)$ bits), but not an index within a $2^{l}$ range. The current window number takes $O(\log n)$ bits, but can be shared across all hash functions. In this case, from the fact that we are currently executing iteration $l$, and the index of the current window, we will be able to derive a range which $i$ may be in. One can verify that this is a range $I$ of $2^{l}$ windows.
@@ 3.2 Concentration bounds
A theorem due to Dubhashi and Ranjan [13, Proposition 5] states that Chernoff-Hoeffding bounds can be applied on negatively associated random variables. In particular, we have the following.
@@
The following proposition from [13, Proposition 7] states some useful properties of negatively associated random variables.
@@
For $j\in[n]$, let $B_{j}$ be the indicator random variable for the event that ball numbered $j$ is sampled. Dubhashi, Priebe, and Ranjan proved in [12, Corollary 11] that the random variables $(B_{j}:j\in[n])$ are negatively associated. It follows from Item 1 in Proposition 3.12 that $(B_{j}:i+1\leq j\leq i+\ell)$ are also negatively associated. Since $X_{\ell,i}=\sum_{j=i+1}^{i+\ell}B_{j}$, $\mathbb{E}[X_{\ell,j}]=\ell k/n$, and the variables $B_{j}$ are identically distributed, the desired bound follows from applying Theorem 3.10. ∎
@@ 4 Streaming algorithm for detecting $\ell_{2}$ super heavy hitter
Let $\mathcal{S}$ be an input stream divided into $L$ windows, each of length $W$. Let $m=LW$ denote the total length of the stream. Let $U$ be a universe of size at most $\mathrm{poly}\,n$ [^6] Every element in the universe can be represented using $O(\log n)$ bits.. $K=(\log n)^{100}$ is a fixed parameter. We say that an element $x$ is a $K$-super heavy hitter if $f_{x}^{2}\geq K\cdot\sum_{y\neq x}f_{y}^{2}$, or equivalently it is a $\left(1-\frac{1}{K+1}\right)$-heavy hitter. We are given a subset of the universe $V\subset U$, and we focus on the substream consisting of elements in $V$. The positive case is that there exists a $K$-super heavy hitter $x^{*}\in V$ (with respect to the substream restricted to $V$) of frequency $L$ and evenly distributed in the original stream. In this section, we develop algorithms that return YES if we are in the positive case, with high probability, return NO if we are far from it (i.e., every element in $V$ has frequency at most $K^{-1/8}L$), and with further guarantees for the intermediate cases. In the following section, we first give a streaming algorithm for the case when $L\geq K^{3/16}$, and in the end, we handle the case when $L<K^{3/16}$ separately.
@@ 5 Streaming algorithm for detecting $\ell_{2}$ heavy hitter
(5.10) $\leq\sum_{v}O\left(\left((K^{-3/4}2^{-l}W\log^{4}n+\log n)2^{l}q\right)^{2}\cdot q\log n\cdot F_{I,v}\right)$